JoelTheBoss (joel_the_boss):

A grid with 3 rows and 52 columns is tiled with 78 identical 2x1 dominoes. How many ways can this be done such that exactly two the dominoes are vertical? @dan815

JoelTheBoss (joel_the_boss):

Details: The dominos will cover the entire board. They are not allowed to jut over the board, or over lap with each other. If I know how to draw it I would... -_-

OpenStudy (dan815):



OpenStudy (dan815):

hmm let ssee

OpenStudy (dan815):

interesting qn!

OpenStudy (dan815):

there are 78 tiles 2 by 1, 78*2=140+16=156 places 52*3=156 tiles

OpenStudy (dan815):

2 tiles are vertical only

OpenStudy (dan815):

so u can repeat this whole thing

OpenStudy (dan815):

like this i think

OpenStudy (dan815):

52 is even so

OpenStudy (dan815):